8÷4(3-1)=?

8÷4(3-1)=?

Grade:12th pass

5 Answers

Vikas TU
9 years ago
solve bu BOD MASS rule.
1 is the ansr.
d.kojalrani
9 years ago
yes the answer is 1,8/4(3-1)=4(3-1)=4(2)=8,8/8=1
TEJKIRAT SINGH
9 years ago
6 is the ansswer.
TEJKIRAT SINGH
9 years ago
srry4 hai ansswer
SRUTI
9 years ago
then th equestion will be like this yaar

(8÷4)(3-1)=? then the answer is 2*2=4